OpenStudy (aravindg):

the quadratic 8 sec^2x-6secx+1=0 has ............ roots .....i need stepwise answer

OpenStudy (anonymous):

replace sec with a variable like a and solve like any other quadratic

myininaya (myininaya):

first solve this for u 8u^2-6u+1=0

OpenStudy (aravindg):

i got 1/4 and 1/2 now what???

myininaya (myininaya):

ok we replaced secx with u since u=1/4 or u=1/2 then secx=1/4 or secx=1/2

myininaya (myininaya):

solve these for x and you are done

OpenStudy (aravindg):

so what is final answer?? how many roots r thr???

myininaya (myininaya):

if it helps you can write in terms of cosine \[secx=\frac{1}{cosx}=> \frac{1}{cosx}=\frac{1}{4}, \frac{1}{cosx}=\frac{1}{2}=> cosx=4, cosx=2\]

OpenStudy (anonymous):

but it cant be possible

myininaya (myininaya):

right

myininaya (myininaya):

becaise cosx is btw -1 and 1

OpenStudy (aravindg):

0 roots???

OpenStudy (anonymous):

-1<cos x<1

OpenStudy (anonymous):

so we have no solution for quadratic equation

myininaya (myininaya):

:)

OpenStudy (aravindg):

no roots u mean???

myininaya (myininaya):

of course

OpenStudy (anonymous):

yes

OpenStudy (aravindg):

wow thx i giv medal

OpenStudy (aravindg):

to both of u

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

:)

OpenStudy (aravindg):

:) more qns coming up..............

OpenStudy (anonymous):

yes:)

myininaya (myininaya):

:)